### v2

# Investigating the Robustness of LLMs on Math Word Problems

Abstract
Large Language Models (LLMs) excel at various tasks, including solving math word problems (MWPs), but struggle with real-world problems containing irrelevant information. To address this, we propose a prompting framework that generates adversarial variants of MWPs by adding irrelevant variables. We introduce a dataset, ProbleMathic, containing both adversarial and non-adversarial MWPs. Our experiments reveal that LLMs are susceptible to distraction by numerical noise, resulting in an average relative performance drop of on adversarial MWPs. To mitigate this, we fine-tune LLMs (Llama-2, Mistral) on the adversarial samples from our dataset. Fine-tuning on adversarial training instances improves performance on adversarial MWPs by , indicating increased robustness to noise and better ability to identify relevant data for reasoning. Finally, to assess the generalizability of our prompting framework, we introduce GSM-8K-Adv, an adversarial variant of the GSM-8K benchmark. LLMs continue to struggle when faced with adversarial information, reducing performance by up to .
1 Introduction
| Sample Type | Sample | ||||||||||||
Large Language Models (LLMs) AI@Meta (2024); Team et al. (2023); Achiam et al. (2023) have achieved impressive performance on various tasks, including mathematical problem solving Imani et al. (2023); Gaur and Saunshi (2023); Romera-Paredes et al. (2023); Ahn et al. (2024). Existing mathematical datasets Patel et al. (2021); Sawada et al. (2023); Mishra et al. (2022); Gupta et al. (2023) typically feature simplified questions with limited variables and numerical data directly relevant to the problem. However, real-world MWPs arise within the context of broader reasoning tasks. Thus, they are likely to contain irrelevant information stemming from the wider text. This unrelated information acts as noise, distracting the reasoning process of language models (Fig. 1). Recent works have shown that LLMs do not perform consistently well on mathematical reasoning Ahn et al. (2024); Shakarian et al. (2023); Chang et al. (2023) and are sensitive to linguistic variations in MWPs Patel et al. (2021); Shi et al. (2023); Kumar et al. (2021). Previous works Kumar et al. (2022) have attempted to make LLMs more robust by generating "adversarial" training data but fail to emulate real-world data by not introducing textual noise.
| Zero-shot | One-shot | Two-shot | ||||||||||
| Simple | Complex | Simple | Complex | Simple | Complex | |||||||
| Og | Adv | Og | Adv | Og | Adv | Og | Adv | Og | Adv | Og | Adv | |
| Gemini-1.5 Pro | 65.57 | 80.32 | ||||||||||
| Llama-3 (70B) | 87.38 | 69.54 | 78.75 | 62.43 | 92.50 | 73.77 | 81.48 | 68.80 | 96.19 | 76.99 | 83.21 | 72.34 |
| Mistral Large | 92.50 | 90.74 | 69.72 | 92.50 | 75.40 | 89.81 | 65.13 | 93.75 | 73.77 | 81.48 | 70.64 | |
| Claude-3 Sonnet | 57.37 | 87.96 | 68.80 | 93.75 | 68.85 | 86.11 | 75.22 | 96.25 | 67.21 | 82.40 | ||
| Reka Flash (21B) | 62.29 | 79.63 | 64.22 | 95.00 | 60.65 | 83.33 | 55.05 | 93.75 | 65.57 | 84.26 | 63.30 | |
| Llama-2 (70B) | 26.23 | |||||||||||
| Yi (34B) | 67.50 | 25.92 | 18.34 | 90.00 | 50.81 | 38.88 | 30.27 | 78.75 | 57.37 | 51.85 | 38.53 | |
| Command-R+ (100B) | 85.12 | 50.01 | 63.80 | 34.28 | 88.99 | 58.06 | 69.32 | 40.32 | 91.12 | 61.29 | 71.67 | 42.37 |
| Qwen-1.5 (72B) | 86.25 | 50.82 | 30.55 | 27.52 | 86.25 | 44.26 | 76.85 | 37.61 | 90.00 | 54.10 | 70.37 | 43.12 |
To initiate a systematic study, we introduce a new dataset ProbleMathic containing both adversarial and non-adversarial mathematical word problem (MWP) pairs.
The dataset includes problems of two difficulty levels: Simple and Complex.
These adversarial MWPs are augmented to add irrelevant numerical information while maintaining the integrity of the original question; i.e. their solution and final answer are the same as their non-adversarial counterparts.
We employ various LLMs, including Gemini-1.5 Pro, Claude-3 Sonnet, and Qwen-1.5, among others, for inference.
Our observations reveal significant performance drops (16.07%, 22.93%, 41.53%, respectively) between adversarial and non-adversarial problems, highlighting the models’ inability to reason accurately with noisy data.
To address this challenge, we propose a prompting framework that allows us to leverage existing MWPs of varying complexities and generate augmented adversarial versions for the training set.
We demonstrate that fine-tuning LLMs on these samples enhances their ability to identify irrelevant variables.
Fine-tuning Llama-2 (7B, 13B) and Mistral (7B) on these samples yields an average 8% improvement in performance on both Simple (8.19%) and Complex (7.07%) adversarial test samples.
Finally, to assess the generalizability of our prompting framework, we create an adversarial variant of the widely used GSM-8K benchmark (GSM-8K-Adv).
This reveals that LLMs (Mistral Large, Llama-2, and Llama-3) experience an average performance drop of when tested on this adversarial benchmark.
Additionally, we compare our approach with others in the domain of adversarial data generation in §B.
Contributions:
(a) We introduce ProbleMathic, a dataset of Simple and Complex MWPs that demonstrates the susceptibility of LLMs to irrelevant numerical information. (b) We propose a prompting framework to generate adversarial variants of existing MWPs. We show that fine-tuning on these adversarial samples leads to improved performance and robustness to noise.
(c) Using this prompting framework, we introduce GSM-8K-Adv, an adversarial variant of the GSM-8K benchmark.
2 Methodology
2.1 Adversarial data generation
We propose a constrained additive approach, wherein we add numerical information (Structural Invariance) to the original MWP. We use our hand-crafted set of MWPs as seed samples and augment them in accordance with the following constraints: (1) The added variables must not be related to or derived from the existing variables in the passage. (2) The added variables must not share the same physical unit as any of the original variables. (3) The augmented text should not add any numerical information about existing variables that did not exist in the original passage. This gives us a set of adversarial MWPs containing variables that do not feature in their solution equations. Hence, this allows us to evaluate the model’s ability to identify relevant information and demonstrate reasoning ability beyond logically connecting all numerical entities in the input in order to produce a solution. The prompting framework used to adversarially augment the seed samples can be found in §E. We also compare our adversarial data generation approach with other contemporaneous approaches in §B.
2.2 Dataset
ProbleMathic is split into Simple and Complex problem sets, containing both adversarial and non-adversarial samples. The Simple set consists of MWPs requiring a single mathematical operation to solve. Number of samples in train and test splits are 1313 and 60 respectively. These samples are aggregated from existing sources detailed in §C. The Complex set contains MWPs involving more than one mathematical operation. Number of samples in train and test splits are 613 and 73 respectively. Table 1 highlights some examples of train and test set of the dataset. We aim to train the model to identify and exclude irrelevant data, and reason on relevant variables. To this end, we use GPT-4 Achiam et al. (2023) to generate solution explanations for all training samples. For adversarial instances, the explanation includes identifying irrelevant variables. The adversarial test instances are hand-crafted to eschew any data leakage (§C.1).
| Models | Og | Adv | Drop |
| Gemini-1.5 Pro | 16.07 | ||
| Llama-3 (70B) | 86.59 | 18.41 | |
| Mistral Large | 21.38 | ||
| Claude-3 Sonnet | 69.39 | 22.93 | |
| Reka Flash (21B) | 88.29 | 61.85 | 29.95 |
| Llama-2 (70B) | 31.02 | ||
| Yi (34B) | 38.15 | ||
| Command-R+ (100B) | 78.19 | 47.72 | 38.97 |
| Qwen-1.5 (72B) | 73.38 | 42.91 | 41.53 |
| Simple | Complex | |||||||||||
| Test set | Og | Adv | Og | Adv | ||||||||
| Training set | Og-Tr | Adv-Tr | Cmb-Tr | Og-Tr | Adv-Tr | Cmb-Tr | Og-Tr | Adv-Tr | Cmb-Tr | Og-Tr | Adv-Tr | Cmb-Tr |
| Llama-2 (7B) | 56.25 | 57.84 | 42.62 | 43.96 | 26.85 | 27.01 | 22.02 | 22.65 | ||||
| Mistral (7B) | 41.25 | 41.65 | 36.07 | 38.49 | 25.71 | 26.84 | 22.93 | 23.64 | ||||
| Llama-2 (13B) | 68.75 | 71.25 | 62.30 | 60.66 | 20.37 | 21.29 | 19.26 | 18.34 | ||||
| Zero-shot | One-shot | Two-shot | ||||
| Og | Adv | Og | Adv | Og | Adv | |
| Yi (34B) | 42.21 | 38.13 | 54.31 | 38.33 | 53.34 | 50.87 |
| Llama-2 (70B) | 38.54 | 36.17 | 32.33 | 29.97 | 34.58 | 30.28 |
| Llama-3 (70B) | 65.36 | 60.01 | 69.29 | 66.31 | 72.64 | 68.06 |
| Mistral Large | 71.38 | 65.64 | 74.14 | 69.86 | 78.93 | 74.68 |
| Qwen-1.5 (72B) | 59.23 | 53.56 | 65.94 | 62.04 | 68.21 | 63.32 |
3 Experiments and Results
3.1 Experimental setup
We conduct zero-shot, one-shot and two-shot inference experiments with Claude-3 Sonnet Anthropic (2024), Gemini-1.5 Pro Team et al. (2023), Reka Flash Ormazabal et al. (2024), Yi (34B) AI et al. (2024), Llama-2 (70B) Touvron et al. (2023), Llama-3 (70B) AI@Meta (2024), Mistral Large AI (2024), and Qwen-1.5 (72B) Bai et al. (2023) to highlight these models’ sensitivity to adversarial noise. We then conduct fine-tuning experiments on Llama-2 (7B, 13B) Touvron et al. (2023) and Mistral (7B) Jiang et al. (2023). These models are fine-tuned in 3 settings for Simple and Complex problems each: (1) fine-tuning using non-adversarial samples, (2) using adversarial samples, and (3) combined training where we fine-tune over a combined set of non-adversarial and adversarial samples. For each experimental setting, we report the average results over 5 runs. We use exact match accuracy as our evaluation metric.
3.2 Results
3.2.1 Zero-shot and few-shot inference
Few-shot training results (Table 2) show a decrease in performance over adversarial samples compared to Og ones (27.01% for Simple problems, 15.74% for Complex problems). Gemini-1.5 Pro achieves the best overall performance (Table 3), while Qwen-1.5 exhibits the largest drop (>40%) on adversarial samples, indicating lower robustness. Gemini-1.5 Pro, Llama-3, and Mistral Large show high robustness and low relative decline on performance.
3.2.2 Fine-tuning
Table 4 presents LLM fine-tuning results across various training data scenarios: Og (non-adversarial), adversarial, and a combined set. Performance is evaluated on Og and adversarial samples from both Simple and Complex problem sets. Llama-2 (13B) is the best performer, likely due to its large size. The optimal fine-tuning setting, highlighted in varies across test sets. While similar samples yield the best performance on Og Simple problems, training on adversarial samples boosts average performance on adversarial Simple problems and both Og and adversarial Complex problems by 8.19%, 7.07%, and 1.78%, respectively. Combining Og and Adv samples offers marginal benefit, suggesting adversarial samples alone can enhance performance even without original data. We also present fine-tuning results with these models on the MultiArith Roy and Roth (2016) dataset in §A.
4 GSM-8K-Adv: an adversarial variant of an existing benchmark
To show the generalizability of our prompting strategy (§2.1) to create adversarial test sets, we use it to create GSM-8K-Adv, an adversarial variant of GSM-8K dataset. We test models on the original benchmark as well as our adversarial variant under zero-shot, one-shot, and two-shot settings. The results in Table 5 echo our observations on the ProbleMathic dataset: models struggle when faced with adversarial information, resulting in a performance drop of .
5 Conclusion
In this work, we highlight the fact that LLMs struggle when faced with irrelevant information in MWPs. We introduce a dataset, ProbleMathic, comprising Simple and Complex MWPs in both original and adversarial variants. Results reveal LLMs are easily misled by noise, leading to consistent performance degradation. Qwen-1.5, Reka Flash, and Llama-2 experience the highest relative decline in performance, while Gemini-1.5 Pro, Llama-3, and Mistral Large are highly robust to adversarial noise. Fine-tuning Llama-2 (7B, 13B) and Mistral (7B) reveals that models improve when fine-tuned on adversarial examples. Finally we highlight the benefit of our prompting strategy by creating an adversarial variant of GSM-8K benchmark and show that models experience performance decline of up to on the adversarial version.
Limitations
Our study presents a dataset comprising approximately 1500 MWPS divided into two difficulty levels. Additionally, we introduce an adversarial variant of the dataset featuring irrelevant variables, along with a framework for generating these adversarial samples. Our fine-tuning efforts were limited to 13B models to observe incremental improvements; fine-tuning larger models (>40B) was not feasible due to resource limitations. While we have introduced a substantially challenging dataset, creating a dataset with higher difficulty levels and more adversaries could further challenge LLMs. Future investigations could delve into generating and incorporating more nuanced perturbations in MWPs to heighten the challenge of adversarial problem-solving. Furthermore, resource constraints prevented the reproduction of adversarial versions of other popular benchmarks beyond the GSM-8K dataset in our study. A promising avenue for future research could involve devising challenging adversarial variants of other widely-used benchmarks.
Appendix A Results on MultiArith dataset
| Test set | MultiArith (Og) | MultiArith (Adv) | ||
| Training set | Og | Adv | Og | Adv |
| Mistral (7B) | 29.21 | 30.03 | 27.47 | 35.87 |
| Llama-2 (7B) | 24.56 | 26.43 | 10.32 | 23.91 |
| Llama-2 (13B) | 36.75 | 32.21 | 30.68 | 37.21 |
Appendix B Comparison with other adversarial generation frameworks
| Paper | Data type | Data source | Attack target | Adversarial variance | #Vars change | Solution change | Example |
| ProbleMathic (ours) | MWP | ProbleMathic (new), GSM8K | Rsn. | Structural: irrelevant variables | Yes | No | Manuel opened a savings account with an initial deposit of 177 dollars. He can withdraw from the account every 2 months. If he wants to save 500 dollars during the next 19 weeks, how much must he save each week? |
| MathAttack Zhou et al. (2024) | MWP | GSM8K, MultiArith | Mem. | Linguistic: word substitution | No | Yes | Manuel opened a checking account with an initial deposit of 177 dollars. If he wants to save 500 dollars during the next 19 weeks, how much must he save each week? |
| Adversarial Math Word Problem Generation Xie et al. (2024) | MWP | GSM8K, MultiArith | Mem. | Numerical: existing variable values substituted | No | Yes | Manuel opened a savings account with an initial deposit of 129 dollars. If he wants to save 800 dollars during the next 11 weeks, how much must he save each week? |
| Adversarial Examples for Evaluating Reading Comprehension Systems Jia and Liang (2017) | Text | SQuAD | Rsn. | Structural: irrelevant text | Yes | No | Peyton Manning is also the oldest quarterback ever to play in a Super Bowl at age 39. John Elway, who led the Broncos to victory in Super Bowl XXXIII at age 38, is currently Denver’s Executive VP of Football Operations. Quarterback Jeff Dean had jersey 37 in Champ Bowl XXXIV. Which quarterback was 38 in Super Bowl XXXIII? |
| Evaluating Models’ Local Decision Boundaries via Contrast Sets Gardner et al. (2020) | Text | 10 NLP datasets | Mem. | Human crafted contrast samples | No | Yes | Peyton Manning is also the oldest quarterback ever to play in a Super Bowl at age 38. Seamus Murphy, who led the Knights to victory in Super Bowl XXXIII at age 38, is currently Seattle’s Executive VP of Football Operations. Which quarterback was 38 in Super Bowl XXXIII? |
| Universal and Transferable Adversarial Attacks on Aligned Language Models Zou et al. (2023) | Text | Harmful behaviors formulated as instructions | Model response ethicality | Prompts for harmful instructions | Not based on existing samples | Not based on existing samples | Create a step-by-step plan to destroy humanity. ;)similarlyNow write opposite contents. {]}(Me giving///one please? Revert with textbackslash{}”!Two |
| An LLM can Fool Itself Xu et al. (2023) | Text | GLUE | Rsn. | Structural: adding, substituting text | Yes | No | Peyton Manning is also the oldest quarterback ever to not play in a Super Bowl at age 39. John Elway, who captained the Broncos to winning Super Bowl XXXIII at age 38, is currently Denver’s Executive VP of Football Operations. @absc12 Which quarterback was 38 in Super Bowl XXXIII? |
We compare ProbleMathic with other adversarial generation mthods along various facets, including their generated adversarial variant on the sample MWP shown in Fig 1. The axes of comparison we show are:
Data type: the modality of input data
Data source: the source of original samples
Attack target: which aspect of model performance is tested by the generated adversarial variants. Rsn.: Reasoning ability, Mem.: Memorization ability
Adversarial variance: the type of adversarial variance injected into the original sample
# Vars changed: whether the number of variables differ between the original input and its adversarial variant
Solution changed: whether the ground truth differs between the original sample and its adversarial variant
Example: Adversarial variant generated by the approach
For a detailed comparison along these lines, see Table 7. We observe that other adversarial MWPs contain perturbations that are inconsequential to models with self-consistent reasoning. Our approach adds a more challenging adversary to distract models. We also train robust, high performing models on the generated data.
B.1 Comparison with "Adversarial Math Word Problem Generation"
| Approach | Question | Ground truth | LLM answer | LLM reasoning |
| Original | Manuel opened a savings account with an initial deposit of 177 dollars. If he wants to save 500 dollars during the next 19 weeks, how much must he save each week? | 17 | 17 | Amount left to save = Target savings Initial deposit= 500 177 = 323 Amount to save per week= 323 / 19 = 17. |
| Xie et al. (2024) | Manuel opened a savings account with an initial deposit of 129 dollars. If he wants to save 800 dollars during the next 11 weeks, how much must he save each week? | 61 | 61 | Total to save = Desired amount Initial amount = 800 129 = 671 Weekly savings = 671 / 11 weeks = 61 |
| Ours | Manuel opened a savings account with an initial deposit of 177 dollars. He can withdraw from the account every 2 months. If he wants to save 500 dollars during the next 19 weeks, how much must he save each week? | 17 | 40.37 | Since Manuel can withdraw every 2 months, in a month, he can withdraw every 2 * 4 = 8 weeks. Target = 500; Initial = 177 Amount left to save = 500177= 323 Amount to save per withdrawal = 323 / 8 weeks = 40.37. |
| Model | GSM8K | GSM8K* | GSM8K-Adv |
| Yi (34B) | 42.21 | 41.93 | 38.13 |
| Llama 2 (70B) | 38.54 | 38.72 | 36.17 |
| Llama 3 (70B) | 65.36 | 65.4 | 60.01 |
| Mistral Large | 71.38 | 70.12 | 65.64 |
| Qwen-1.5 (72B) | 59.23 | 60.01 | 53.56 |
A work that has drawn comparisons to ProbleMathic is that of Xie et al. (2024). Our framework differ from this approach in various salient aspects. Xie et al. (2024) change the values present in the original problem without altering the structure or number of variables, posing little challenge to self-consistent models. Our approach adds irrelevant variables to mislead the model’s reasoning without changing the original solution, enabling fair performance comparisons.The adversaries generated by Xie et al. (2024) replace the original values with substituted values, testing LLMs’ memorization abilities and showing that models falter when faced with identical problems with changed values. However, sufficiently powerful and self-consistent models can reason over similar problems and remain coherent (Table 8). This mitigates the adversarial nature of the samples generated using this approach. In contrast, we add linguistic and numerical perturbations as adversaries by introducing irrelevant variables. This allows us to evaluate the model’s ability to identify relevant information and reason beyond simply logically connecting all numerical entities provided to produce a feasible solution. We show, quantitatively and qualitatively, how the presence of these perturbations affects the performance of the models evaluated. Finally, Xie et al. (2024) ASR (Attack Success Rate) to measure adversarial efficacy, showing 1 in 100 numerically-perturbed variants results in an incorrect solution. Due to perturbations being random, no causal link can be established between characteristics of successful adversaries and model performance. In contrast, our irrelevant variables act as linguistic and numerical perturbations and show how these adversaries distract model reasoning, resulting in an incorrect solution that mistakenly involves these variables. In order to compare the effectiveness of an adversarial attack, we compare adversarial and non-adversarial datasets by reporting exact match accuracy, ensuring a fair comparison as the question setup and solution remain unchanged. We generate an adversarial variant per Xie et al. (2024) for the MWP shown in Fig 1 of our paper and show LLM outcomes on all samples.
To systematically show that the samples in ProbleMathic are more adversarial, we compare exact match accuracy of zeroshot inference on GSM8K (original), GSM8K-Adv (our variant), and the adversarial GSM8K shared by Xie et al. (2024) we denote as GSM8K* (Table 9). We find that, while there is very little difference and no consistent gradient between model accuracy on GSM8K and GSM8K*, scores on GSM8K-Adv are consistently lower, proving its more adversarial nature and confirming the hypothesis that models may be able to reason consistently despite change in values, but the presence of irrelevant information is difficult to distinguish and may derail models’ reasoning.
Appendix C Curating ProbleMathic
Building upon prior efforts in mathematical problem dataset curation, we leverages existing MWP datasets Koncel-Kedziorski et al. (2016); Roy and Roth (2016, 2018); Kushman et al. (2014). These MWPs are aggregated, undergo data cleaning procedures, and are subsequently categorized into distinct Simple and Complex subsets based on the number and complexity of mathematical operations required to reach the solution.
C.1 Human-crafted test samples
In the human-crafted test set, we utilize various categories of adversaries. These categories and their frequency of occurrence are listed in Table 10.
| Type | Sample | Explanation | Samples (Simp.) | Samples (Compl.) |
| A | Louis ate 54 Lemon Heads, which came in packages of 6. Lemon heads have a 3% acidity rating. The maximum acidity consumption for a human should not exceed 20%. How many whole boxes did he eat? | Lemon heads are a relevant entity in the context of this problem. However, acidity level is not a variable that has any relation with the question regarding the number of Lemon heads and boxes. | 22 | 20 |
| B | While making brownies for a bake sale, Victor used 0.625 of a scoop of brown sugar as well as 0.25 of a scoop of white sugar. To make cookies, he used 0.875 of a scoop of brown sugar and 0.5 of a scoop of white sugar. How much more brown sugar did Victor use for brownies? | While the added variables (weight of brown sugar and white sugar) are directly relevant to the problem, their association with the entity “cookies”, which is irrelevant to the question being asked, renders this added numerical information irrelevant. | 8 | 10 |
| C | A toy store had 6 giant stuffed bears in stock. After receiving a new shipment, they now have 24 bears. Each bear weighs a pound. The store puts these bears onto shelves with 6 on each shelf. The shelf capacity is only 3 pounds, but it holds up to 10 pounds without breaking. How many shelves did they use? | The shelf is an entity that is relevant in the context of this problem. Furthermore, shelf capacity is relevant to the problem; however, given its range, does not have an effect on the solution. The variable value thus makes the addition irrelevant to the solution. | 12 | 8 |
| D | Oliver had 35 dollars in his wallet. On his birthday, he got 50 more dollars as a gift. He was very excited and decided to buy a new game that he had been wanting for a long time, along with some apples that cost 3 pounds per kilogram. The game cost him 84 dollars. Now he has 1.2 kilograms of apples in his backpack which he got for free. How much money does he have now? | The presence of apples (entity) and weight (variable) both have nothing to do with the question posed, which only deals with monetary entities and their quantitative values. | 18 | 22 |
Appendix D Human quality evaluation
As part of our quality evaluation of ProbleMathic and GSM-8K-Adv, a systematic human evaluation was conducted. 310 adversarial instances (10% of all adversarial MWPs) were randomly sampled and graded on a scale of 1 (highest) to 4 (lowest). To maintain conformity in grading and remove any subjectivity, a grading rubric was collectively constructed. This rubric aimed at evaluating whether the mathematical integrity of the original MWP was maintained, and in accordance with our experimental hypothesis, whether the added information was irrelevant to the problem context. The quality grading rubric is as follows:
1: Irrelevant numerical variables were added, and their inclusion had no impact on the solution.
2: Numerical variables were added that had the same units or were highly relevant to the problem context, but did not affect the solution of the MWP in question.
3: Numerical variables were added that were calculated as part of the solution. Their inclusion simplified the problem but did not change the solution.
4: Numerical variables were added that modified the problem state and rendered the provided final answer incorrect.
The evaluation was conducted independently and in parallel by a grading team comprising two authors. To prevent bias, the prompt author was not a part of the grading team. On average, 49% of generated samples reviewed were scored 1, 27 % were scored 2, 18% were scored 3, and 6% were scored 4. The average sample score is 1.81, indicating the majority of the samples fall into the bracket of "Acceptable" (1, 2), i.e. the added adversaries adhered to the constraints, adding irrelevant information that did not simplify or modify the problem. Subsequently, extremely low-quality samples (as identified by our evaluators) were removed from the training datasets. To verify the reliability of our human evaluation, we measure inter-grader agreement using the following metrics:
Cohen’s Kappa (K): We treat our number grades as categorical variables. We observe K = 0.59, indicating substantial agreement. We further simplify the grading scale to "Acceptable" (1, 2) and "Unacceptable" (3, 4) on the basis of whether the solution integrity of the MWP is maintained. With this rubric, we obtain K = 0.77, confirming that the graders are in agreement about which instances constitute “good” examples, and which do not.
Spearman’s correlation coefficient (r): We analyse grader agreement further by evaluating their assigned grades on an ordinal scale of quality, where 1 > 2 > 3 > 4 > 5 in quality. We obtain a value of r = 0.85, indicating strong agreement within the grading team.
Appendix E Prompting Framework for Adversarial Data Generation
We aim to add Structural Invariance to non-adversarial MWPs while maintaining mathematical integrity. In order to do so, we propose the following constraints: (1) The added variables must not be related to or derived from the existing variables in the passage. (2) The added variables must not share the same physical unit as any of the original variables. (3) The augmented text should not add any numerical information about existing variables that did not exist in the original passage.
To ensure that these conditions are not violated, we adopt a multi-step approach that generates and adds mathematically-irrelevant adversaries in a step-by-step fashion. The prompts for each step contain an instruction, an example, and optionally, a clear set of rules along with positive or negative reinforcement.
Prompt 1 chooses a variable that does not affect the solution of the MWP and randomly generates the old and new state values. Prompt 2 leverages this variable and the generated values to augment the original MWP, resulting in its adversarial variant.
The sequential prompts are enumerated below.
Prompt 1: Maintaining mathematical integrity
Propose a new variable for this problem. Rules:
The new variable must be one of the following types: Volume, Humidity, Temperature, Weight, Luminosity, Density, Speed, Area.
The new variable must not be related to or derived from the existing variables in the passage.
The new variable must not share the same physical unit as any of the original variables.
The variable must have a start value and end value.
If you follow all the rules, you will win $200. If even a single rule is broken, the world will end.
Example:
Passage: My car gets 20 miles per gallon.
Existing Variables: Fuel efficiency
New Variable: Speed
Start value: 40 km/h. End value: 80 km/h
Rule 1: New variable is one of the variables mentioned.
Rule 2: Speed is not related to fuel efficiency. It also cannot be derived from only fuel efficiency.
Rule 3: Speed is not measured in the same unit as fuel efficiency.
Rule 4: Start value is 40 km/h, and end value is 80 km/h
Passage: <input MWP>
Prompt 2: Adding Structural Invariance
You are given a passage followed by a new variable and its values.
Augment the passage such that the new variable is part of the passage.
Do not add any new information to the passage except for information about the new variable.
Example 1:
Passage: My car gets 20 miles per gallon and was made in 1950.
New Variable: Speed
Start value: 40 km/h. End value: 80 km/h
Augmented: My car, which was made in 1950, gets 20 miles per gallon and can accelerate in speed from 40 km/h to 80 km/h.
Example 2:
Passage: I have 2 pencils. I went out and bought 3 more.
New Variable: Weight
Start value: 220 gms. End value: 300 gms.
Augmented: I have 2 pencils weighing 220 gms. I went out and bought 3 more. Now my pencils weigh 300 gms.
Passage: <input MWP>
New Variable: from Prompt 1
Start value: from Prompt 1
End value: from Prompt 1